how to write a product using exponets
1
Expert's answer
2014-03-14T07:41:40-0400
It is easier to explain using example.
If you have some value of the same multipliers then you can write it as
a*a*a*...*a (n times) = a^n.
If you have different multipliers then we get
a*a...*a (n times)*b*b...*b (k times) = (a^n)*(b^k).